OpenStudy (anonymous):

Choose the correct slope of the line that passes through the points (1, -3) and (3, -5).

OpenStudy (anonymous):

@chycora

OpenStudy (anonymous):

@Write0The0Right0For0Wrong0Doing

OpenStudy (anonymous):

@nothingwasthesame

OpenStudy (anonymous):

need help

OpenStudy (anonymous):

@prettygirl_love

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ 1}}\quad ,&{\color{blue}{ -3}})\quad &({\color{red}{ 3}}\quad ,&{\color{blue}{ -5}}) \end{array} \\\quad \\ % slope = m slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}}\)

OpenStudy (anonymous):

sup

OpenStudy (nothingwasthesame):

hi

OpenStudy (anonymous):

wyd

OpenStudy (nothingwasthesame):

im bout to help you

OpenStudy (prettygirl_love):

what do u need help with

OpenStudy (jdoe0001):

soup? yes please, with some carrots and zuchinis

OpenStudy (prettygirl_love):

what are you talkin about

OpenStudy (anonymous):

will some one help me

OpenStudy (prettygirl_love):

WITH WHAT!!!!

OpenStudy (anonymous):

the problem the slope

OpenStudy (prettygirl_love):

the what

OpenStudy (anonymous):

(1, -3) and (3, -5).

OpenStudy (anonymous):

the slope

OpenStudy (prettygirl_love):

oh

OpenStudy (prettygirl_love):

sorry can not help with that

OpenStudy (anonymous):

@l_o_w_key

OpenStudy (anonymous):

y - 6 = one half(x + 3) y = one halfx - 6 y + 3 = one half(x - 6) x - 2y = 12

OpenStudy (chycora):

@TheEdwardsFamily

OpenStudy (anonymous):

help me

OpenStudy (triciaal):

the slope is the change in y divided by the corresponding change in x change means difference (you subtract) for 2 points (x1, y1) and (x2, y2) the slope , m m =( y2 -y1) / (x2 - x1) use the points given and calculate see above response by @jdoe0001

OpenStudy (anonymous):

the slope is -1